The hypotenuse of a right triangle is 13 inches. Its perimeter is 30 inches. Find the length of the shortest side.
Answer by sachi(548)

The hypotenuse of a right triangle is 13 inches. Its perimeter is 30 inches. Find the length of the shortest side.
ans:-
in the right triangle hypotenuse h=13 ,p is perpendicular & b is base
so h+p+b=30
or p+b=30-13=17 -----1
p^2+b^2=h^2=13^2=169 ------2
now (p-b)^2+(p+b)^2=2(p^2+b^2)=2*169=338
or (p-b)^2=338-(p+b)^2=338-17^2=338-289=49
or p-b=sqrt49=7----3
now solving 1 & 2 we get p=12 & the shortest one is b=5" ans